OpenStudy (anonymous):

completely simplify 2ln e^(5x)

OpenStudy (anonymous):

log and exp are inverse functions, meaning \(\ln(e^x)=x\)

OpenStudy (anonymous):

so would that be 10x or am i doing that completely wrong?

OpenStudy (anonymous):

yes it would be \(10x\) since \(\ln(e^{5x}) =5x\)

OpenStudy (anonymous):

ok. thanks so much